Vapour density of PCl5 is 104.25 at ToC. Then degree of dissociation of PCl5 is. (Mw = 208.5)

Grade:12th pass

2 Answers

Vikas TU
7 years ago
From the formulae:
1 + a = Mth./Mobs.
1 + a = 208.5/(104.25*2) => 1
a = 0
Hnece degree of dissociation is zero which cannot be posible hence the eqm. would not attain.
Kapil jain
5 years ago
a = (D-d)/(n-1)d 
Here a is alpha()degree of dissociation), D=vapour density before dissociation=M react/2
d=vapour density after dissociation=M obs/2
n =number of moles of product obtained by one mole of reactant
Now here D=208.5/2 = 104.25
And d= 104.25(given)
And n = 2/1=2
On putting theSe values we get a = 0